### v1

# BATIS: Bootstrapping, Autonomous Testing, and Initialization System for Quantum Dot Devices

Abstract
Semiconductor quantum dot (QD) devices have become central to advancements in spin-based quantum computing. As the complexity of QD devices grows, manual tuning becomes increasingly infeasible, necessitating robust and scalable autotuning solutions. Tuning large arrays of QD qubits depends on efficient choices of automated protocols. Here, we introduce a bootstrapping, autonomous testing, and initialization system (BATIS), an automated framework designed to streamline QD device testing and initialization. BATIS navigates high-dimensional gate voltage spaces, automating essential steps such as leakage testing and gate characterization. The current channel formation protocol follows a novel and scalable approach that requires a single measurement regardless of the number of channels. Demonstrated at on a quad-QD Si/SixGe1-x device, BATIS eliminates the need for deep cryogenic environments during initial device diagnostics, significantly enhancing scalability and reducing setup times. By requiring minimal prior knowledge of the device architecture, BATIS represents a platform-agnostic solution, adaptable to various QD systems, which bridges a critical gap in QD autotuning.
I Introduction
Quantum dot (QD) qubit devices are commonly formed by electrostatically manipulating the potential energy landscape of a two-dimensional (2D) electron or hole gas (EG or HG). This procedure produces interleaved regions of accumulated and depleted charge density that serve as QD qubits, charge sensors, and charge carrier reservoirs [1, 2, 3, 4, 5]. In recent years, the number of QDs in such devices has risen rapidly [6, 7, 8]. In the near term, human intervention will no longer be possible to tune up many dots simultaneously. The field of autotuning has arisen to address this problem.
Modern semiconductor QD qubit devices have two critical features that enhance their performance and simultaneously make autotuning challenging and essential. First, the active region of the device—the QDs, charge sensors, and reservoirs—contains no doping. All charge carriers—electrons or holes—are induced by gate voltages. No current can flow until gate voltages are applied and carriers are accumulated, yet those voltages are unknown a priori. Moreover, the exact voltage configuration changes not only between cooldowns but even between successive initialization steps. Second, QD device chips have multiple current channels that often serve different purposes. Some can host an array of QD qubits, while others provide a home for charge sensors. These channels are not separate until pinch-off voltages, i.e., the gate voltage threshold at which carrier accumulation is eliminated, are determined. Moreover, since the accumulation and depletion of carriers must occur in alternating regions separated by only tens of nanometers to enable the formation of QDs separated by tunnel barriers and proximal charge sensors, these channels must interact.
Here, we describe a bootstrapping, autonomous testing, and initialization system (BATIS) that can automatically test the device and set the gate voltages to ensure that the QD chip is properly accumulated and all required pinch-off voltages are determined. BATIS takes as input the device gate layout and a high-level understanding of the desired operation, including critically where current channels should be continuous and which gates should divide one channel from another. The outputs of BATIS are the gate voltage parameters that make the QD chip functional, with all current channels well delineated and all pinch-off voltages known. Importantly, in this work, the sample is kept at around –about 10 times warmer than other typical devices [9, 10]. Such temperatures are highly desirable for what is known as high-temperature operation [11]. However, higher temperature causes an increased difficulty in thresholding pinchoffs as the coulomb blockade and all measurement features derived therein are less well defined and easier to confuse with noise [12]. BATIS is capable of tuning at these high temperatures, which is important both for scalability and automated testing as, for example, is possible using cryogenic wafer probers [13, 14].
Previous work on autotuning has shown that a bootstrap process is sufficient for further tuning of QD devices. For example, if current channels are known in modulation-doped QD devices, the formation of double QDs with few-electron occupation can be automated [15]. In such devices, autonomous bootstrapping has been demonstrated [16, 17]. If a charge sensor is available, undoped SiMOS devices have been autotuned to achieve a single-electron QD [18]. For nanowires, which by design have a single, known current channel, automation into fully formed QDs has been demonstrated, provided device-specific details are placed into a configuration file [19].
For accumulation-mode, undoped QD devices, previous work has enabled progress in tuning device topology [20, 21] and charge-state calibration [16, 22]. Methods for automated detection of unintentional, spurious QDs have also been proposed [23]. More recently, a framework to achieve orthogonal control of the electrostatic potential landscape in a large 2D array has been put forward [24]. However, autonomous bootstrapping for accumulation mode QD devices has never been attempted, and most tuning algorithms proposed to date assume that steps such as device characterization and the definition of the QD(s) channel(s) have already been completed [25].
The approach for designing BATIS differs from others, typically exercised by the community. All bootstrapping algorithms proposed to date are highly device-specific, with the algorithm’s overall efficiency measured as progress-per-second or some other metric against time used as a figure of merit. However, time-based metrics for device-specific algorithms provide little basis for valid comparison against alternative methods. Such metrics quickly become instrument-bound. Taking large numbers of data points per measurement, each with integration times in the or , quickly dwarfs the simplest point-wise threshold analysis performed on modern multi-core CPUs clocked in the . Users can easily speed up their benchmarks for single-use cases by upgrading to faster instrumentation.
Moreover, the resulting algorithms provide single-use, device-specific applications by assuming substantial knowledge about the tuned state of QD devices. Thus, even if such specialized algorithms perform well on the development device, they require significant redesigning to be deployable on other QD devices.
In contrast, when creating BATIS, we focused on developing a platform-agnostic tool that can be applied to a variety of QD devices. While the procedural flow of BATIS is predetermined, the executable flow is automatically adjusted based on the general device architecture and connectivity to the measurement instruments, both provided within the configuration files. In other words, rather than being programmed to follow a fixed sequence of measurements, BATIS is designed to choose which steps to take and change the expected outcome of each step – depending on the device used. Further adjustments are automatically made as the tuning progresses and the algorithm learns the device.
The device characterization and tuning processes are a repetitive yet essential part of initializing QD-based qubit chips. The majority of autotuning efforts have focused on the latter phases of automatic device tuning, including coarse tuning, charge-state tuning, virtualization, and fine-tuning, which is sufficient for quantum dots in a one-dimensional channel or when the initial step is performed by hand [25]. In modern chips, the scale of the tune-up will soon preclude hand-tuning as fully 2D structures become ubiquitous. BATIS addresses the challenges that are about to become limiting in this widespread class of quantum semiconductor chips.
II BATIS Design
Figure 1(a) shows a gate-defined QD device that illustrates the critical features of accumulation-mode, undoped devices: multiple channels (three white arrows), regions for forming quantum dots (four in the qubit channel, below, and two single-QD charge sensors, above), and reservoir gates to connect these regions to ohmic contacts for current flow. This device exemplifies a commonly used fabrication strategy known as overlapping gates [27, 28].
The screening gates, shown in yellow, are the closest layer of gates to the top gate oxide in the three-layer gate design and are used to isolate each of the channels from each other and from other regions in the chip. The second layer (shown in blue and red) includes the reservoir and plunger gates. The reservoirs enable the flow of electrons from the ohmic contacts toward the center of the device where QDs will form in the electrostatic potential profile controlled by each plunger gate. The final layer consists of barrier gates (grey), which control the tunnel couplings between QDs—both qubit QDs and charge-sensing QDs—and between the outermost QDs and reservoirs. By carefully adjusting the voltages applied to each of the gates, the quantum dot qubit chip with charge sensors is made functional [2, 3, 4, 5].
BATIS is designed to automate the task of testing and characterizing all gates and forming all necessary 1D current channels for arbitrary-length linear QD arrays. The procedural flow of the single button press algorithm is depicted in Fig. 1(b).
BATIS takes as input two configuration files describing the device design and certain generic characteristics of the device and instruments. Based on that information, it performs full diagnostics and characterization of each gate and determines the proper configuration for the consecutive stage of the tuning processes. Each of the six stages in BATIS involves at least one device measurement followed by data analysis to test the functionality of a selected subset of gates, shown in Fig. 1(b) by gray rounded rectangles. The optional illumination stage is executed on an as-needed basis. Illumination involves exposing the QD device to 780 light under carefully chosen gate voltages, to modify its state. Such exposure shifts the gate voltage response by modifying the amount (and if desired, the sign) of charge trapped at the semiconductor-oxide interface [26].
The algorithm follows a conditional logic visually represented with arrows of different colors and textures in Fig. 1(b). The solid gray arrows indicate the procedural step in the sequence. Passing a test for a current stage, marked with yellow diamonds in Fig. 1(b), initiates a successive measurement and assessment stage, as indicated with solid green arrows. Failing a test can result in repeated measurements, device illumination, or tuning termination, depending on the stage. The measurement schemes, together with idealized and real-world data associated with each stage, are shown in Fig. 2.
Once completed, BATIS produces a device diagnostics file that stores all information about each gate learned during tuning. It also pre-configures the QD device for further use. The details of the flow of the algorithm are discussed in the following sections.
II.1 Initialization and First Leakage Test
The two configuration files used as input for BATIS include descriptions of all gates, desirable current channels, and the QD device connections. The files include also details about the software configuration necessary to execute BATIS and a list of variables that need to be considered at each stage. These variables consist of instrumentation limits to ensure that the algorithm does not search in ranges not supported by the instruments and device limits to prevent the device from breaking during tune-up. Details included in configuration files allow the linking of abstract concepts about the device components with the language used in BATIS.
The functionality of gates is determined automatically from the device architecture that the end user inputs, building a gate-lookup table. The files also contain a wire map and instrument information detailing the measurement system setup such that BATIS can automatically read and execute all specialized measurements, connecting the gate-lookup table with the measurement instruments. This allows BATIS to remain device- and instrument-agnostic.
BATIS starts by checking the device for leakage to the fridge ground through the oxide. The first leakage measurement happens before the electron accumulation during the leakage test stage, see Fig. 1(b). The resistances of each connection are measured with respect to each other and compiled into a leakage matrix, as shown in the second and third row of Fig. 2(a). The desired result is a high resistance between all the connections, an entirely yellow matrix.
Defects during the fabrication process can result in leakage. Leakage manifests as off-colored cells in the matrix, as shown in the right panel in the second row and in the third row of Fig. 2(a). Should leakage be detected, BATIS terminates and flags the device as unusable.
II.2 Global Accumulation
For accumulation mode QD devices no 2DEG is present at zero applied gate voltage. Applying a non-zero bias lowers the conduction band minimum below the Fermi level, accumulating a 2DEG in the quantum well near the Si/SixGe1-x interface. The goal of the global accumulation stage, see Fig. 1(b), is to determine the characteristic voltage where enough of the 2DEG is accumulated that a significant amount of current begins to flow between the ohmic contacts, effectively turning the device on.
Idealized successful and failing measurements for this stage are visualized in the left and right panels in the middle row of Fig. 2(b), respectively. The turn-on process in real QD devices is not always straightforward, as even a good device might not turn on properly when first tested. If the algorithm encounters a problem within global turn-on, it has a recovery pathway through optional illumination, see Fig. 1(b).
Illumination is initiated when the device fails to turn on or the turn-on voltage is much too high to be usable. After illumination, the global accumulation stage is re-initiated. If the preference is not to use illumination, BATIS allows skipping this step and going straight to repeating the global turn-on stage. Whether or not illumination should be used is an option that is pre-defined in the configuration file.
II.3 Accumulated Leakage Test
Once the 2DEG is accumulated, the second, accumulated leakage test is initiated. This step proceeds analogously to the first leakage check, except that it does not include the ohmic connections. This is because this leakage check is taken when the device is accumulated, and the ohmics are intentionally shorted to each other. If this stage fails, it indicates that a gate is shorted to the 2DEG flowing beneath it, which is an unrecoverable fabrication defect. Should that occur, BATIS terminates.
II.4 Screening and Reservoir Characterization
The turn-on characteristic voltage is taken globally, meaning it does not provide gate-specific information. To learn that information, BATIS performs a series of individual gate characterizations. The first of these is performed at the screening and reservoir characterization stage in Fig. 1(b).
With the 2DEG accumulated and the device turned on, BATIS depletes the voltages on individual gates until the current on the corresponding channel is pinched off, as displayed in the right panel in the second row in Fig. 2(c). BATIS uses these pinch-off curves to determine where to place reservoir gates for proper 2DEG accumulation and establish bounds to operate the screening gates.
The typical failure modes for this stage include too much depletion and too much accumulation. If the channel below a gate is extraordinarily depleted, no voltage can be applied to the gate to accumulate it. However, since the screening and reservoir characterization stage is only initialized if all current channels turn on properly at the global accumulation stage, the inability to accumulate implies either an issue with the measurement instrument or hysteresis with the gate’s action upon the 2DEG. Similarly, pinching off of all current channels is also confirmed at the global accumulation stage, and thus, it is implausible to observe too much accumulation at this stage. To test whether the issue is caused by the instrument, the algorithm executes a repeated measurement. If the error is observed a second time, it indicates an issue with the gate and the algorithm terminates.
II.5 Forming 1D Current Channels
Selectively removing 2DEG from certain regions of the device is done to form the 1D current channels. It requires depleting the screening gates and accumulating the plungers and barriers (often collectively called finger gates) such that the current flows between the screening gates along a desirable path, see the left panel in Fig. 2(c).
Determining voltages where 1D channels are properly formed is typically done by sweeping two screening gates simultaneously in a 2D sweep while holding the finger gates constant. The formation of the channel manifests as a triangular-sloped region (the so-called triangle plots). If the finger gates are not accumulated enough, this region will not form. Thus, multiple scans with incremented finger gate voltages—a three-dimensional dataset—are often required before the region is observed for all required channels.
BATIS employs a different approach to simultaneously accumulate multiple 1D channels using only a single two-dimensional dataset. Instead of sweeping two screening gates, BATIS takes a 2D voltage scan while sweeping one screening gate against all of the finger gates between that gate and the other screening gate used to form the 1D channel (typically the central screening gate). Since BATIS already learned the necessary configuration for the channel isolation with the central screening gate, this information is now utilized to configure the remaining screening gates. The 1D channel formation technique implemented in BATIS requires a single measurement for each pair of gates, which makes it much more efficient than the standard approaches.
The plot in the right panel in the second row Fig. 2(d) exhibits three characteristic regions: a no-current region in the bottom left, a screening-gate-dominated region in the bottom right, and a finger-gate-dominated region in the top left. BATIS automatically selects an operating point in the top left part of the plot, near the boundary between the current and the no-current regions. The automatic analysis of the regions and their respective boundaries is enabled by fitting a synthetic model to the experimental data, similar to the presentation in [29]. Additional criteria based on the local properties, e.g., estimated resistance and derivatives of the measurement, are used to optimize the selection of the final point among numerous candidates in the operating region.
If the finger-gate-dominated region does not form, BATIS illuminates the device with the finger gates at a voltage 400 more negative than the most recent illumination. After such an illumination, the entire device response can change, because of cross-talk and the interdependence of the current paths. Thus, BATIS resumes autotuning from the global accumulation step.
II.6 Finger Gate Characterization
Once isolated current channels are formed, finger gate pinch-offs can be measured. This is done by assuming that only the nearest neighbors can affect the accumulation under a gate via crosstalk when pinching off the channels. Ideally, as the voltage on a given finger gate is lowered, the pinch-off curves should exhibit a smooth, sigmoidal trend, eventually cutting off the current in the respective channels. An idealized representation of such a curve is visualized in the left panel in the second row in Fig. 2(e).
The bottom panel in the third row of Fig. 2(e) shows a typical curve for these measurements. As the voltage is reduced, the curve begins to exhibit the characteristic Coulomb blockade oscillations typical of QD devices [12]. In principle, such oscillations could complicate any fit-based analysis. Instead, BATIS uses thresholding and slope detection to determine the pinch-off voltage on a point-by-point basis, going in the direction of increasing voltage (the reverse of how the signal was measured) to evade the small deviations of background noise, whilst capturing the first significant turn-on.
III Experimental Demonstration
We now demonstrate BATIS on a 1D quadruple Si/SixGe1-x QD device, nominally identical to the one shown in Fig. 1(a). The device’s schematic consisting of 24 gates is shown in Fig. 3(a). The first layer of gates is made up of four screening gates, Si for , shown in Fig. 3(a) in yellow. The next layer consists of five reservoirs, Ri for , and six plunger gates, Pi for . The reservoirs are shown in Fig. 3(a) in blue, while the plungers are shown in red. The final layer consists of nine barrier gates, Bi for , shown in Fig. 3(a) in gray.
The goal of BATIS is to form three current channels in the quad-QD device, as indicated with white arrows in Fig. 3(a). BATIS brings a grounded but not characterized device from Fig. 1(a) to a fully operational device in Fig. 1(d). In what follows, we explain how BATIS performs the quad-QD device bootstrapping, testing, and characterizations.
First Leakage Test. BATIS begins by performing a device leakage test. For the quad-QD device used in our demonstration, the sample mount holding it contains connections, of which are used for instrument channels to control individual gates and ohmics and an additional wired channels are unused for this device. Thus, the size of the leakage matrix is to examine all possible leakage states. BATIS first checks the diagonal elements of the leakage matrix. If the resistance of every connection exceeds 25, BATIS declares that there is no leakage in the device.
If any of the diagonal elements falls below the 25 threshold, the device is declared leaking. Depending on where the current leak occurs, it is sometimes possible to recover a leaking device. Thus, rather than terminating, BATIS continues to fill in the leakage matrix for diagnostic purposes. Due to the symmetry of the measurements, the leakage matrix is also symmetric and only the lower triangular part of the matrix needs to be completed. Moreover, passing the threshold test by a diagonal component automatically guarantees that the entire column corresponding to that component also passes the test. Thus, only the columns corresponding to the elements below 25 need to be determined. BATIS continues to fill the leakage matrix one column at a time. In the worst case, a total of resistance measurements are necessary to complete the leakage matrix.
The leakage test needs to be executed only once per cool-down if under fully automatic control. This is because BATIS has built-in safety checks during voltage choice selection to avoid exceeding safe voltage breakdowns between neighboring gates for all measurements.
Global Accumulation. Once no leakage is confirmed, BATIS performs the global turn-on stage. In our demonstration, the algorithm executed three rounds of illumination before the observed global turn-on values were determined to be satisfactory, occurring between and . During the first two illuminations, all gates were grounded to reset the device after cooldown and verify that no instruments were erratic. After that, for the third illumination, BATIS chose to apply voltages to all screening and reservoir gates. The voltages required to achieve turn-ons at around , calculated following the protocol in Ref. [26], were set to for gates associated with channel I1 and I2, and for gates associated with channel I3. The central screening gate, S2, was left grounded during the illumination.
The fourth iteration of global turn-on measurement was successful. The summary of all four runs is presented in Table 1. The resulting configuration of the QD device is shown in the left panel in Fig. 1(c), where the pinch-offs on all the screening gates, excluding S2, and all of the reservoir gates are colored based on the learned gate voltage.
| Channel | Run 1 | Run 2 | Run 3 | Run 4 | |
| characteristics | |||||
| I1 | VP () | ||||
| Vmax () | |||||
| I2 | VP () | ||||
| Vmax () | |||||
| I3 | VP () | ||||
| Vmax () | |||||
Accumulated Leakage Test. With the global accumulation confirmed, BATIS proceeds to the accumulated leakage test. A second leakage matrix is determined, following the same strategy as during the first leakage test. Like previously, BATIS finds that all resistances exceed 25 . Since the ohmic connections are excluded, the resulting matrix has five fewer indexes. As none of the connections leaked, only measurements following the matrix diagonal were taken.
Screening and Reservoir Characterization. The next stage in BATIS involves the characterization of the screening and reservoir gates. The results of these characterizations for all tested gates are presented in Fig. 3(b). Unlike the other screening and reservoir gates, S2 needs to be measured simultaneously on channels I1, I2, and I3 since it borders all of the current channels. Because of this, the data for the S2 pinch-off can be found in Fig. 3(c) as it needs to be organized by channel.
| Gate | VP () | VI () | VO () |
| S2 | |||
| S1 | |||
| R1 | |||
| R2 | |||
| S3 | |||
| R3 | |||
| R4 | |||
| S4 | |||
| R4 | |||
| R5 |
The voltage values are quite consistent between the sets of reservoir and screening gates intended to form the desired current channels, a feature from the illumination from earlier [26]. We note that for gate S1, BATIS acquired signal over , i.e., twice as long as for all remaining gates. This is likely due to poor fit scoring using the reduced voltage scan range, which caused BATIS to perform additional measurements with an increased sweep range to be more confident in the derived pinch-off value. Since the long tail of the signal has no features, the signal shown in Fig. 3(b) is truncated to be consistent with the remaining measurements for visualization purposes.
BATIS sets the reservoir gates two standard deviations above the mean of the phenomenological fit it performed on the signal’s gradient. This is to ensure that reservoirs will consistently provide electron flow through the channel as we do not want the cross-talk effects to cause drastic changes in the electron accumulation. At the same time, if the reservoirs are too accumulated, they can form spurious channels behind the regions of interest. To resolve this, the screening gates are set below their pinch-offs in Fig. 3(b) to prevent current flow through spurious channels 111The derivative of the screening gate pinch-off was fit to a function like . The non-central screening gates were set to be before the mean of the derivative..
The summary of the pinch-off voltages and the operating point selected by BATIS is presented in Table 2. The resulting configuration, where the device is in the correct operating point to perform channel accumulation, is visualized in the central plot in Fig. 1(c).
Forming 1D Current Channels. With all screening and reservoir gates characterized, BATIS proceeds to form the 1D current channels. The quad-QD device requires three current channels to be formed, as indicated with the white arrows in Fig. 3(a). BATIS identified three different operating points, one for each 1D current channel calculated from the simultaneous channel accumulation in Fig. 3(c). The summary of the screening and finger gates configuration for each operating point is presented in Table 3, and a collapsed zero cross-talk visualization is pictured in the right panel in Fig. 1(c). The three points are consolidated into one where the voltage color for each gate is the VO indexed by the channel the gate belongs to.
| Gate | V () | V () | V () |
| S1 | |||
| S3 | |||
| S4 | |||
As we discuss in Sec.II.5, the algorithm for 1D current channel formation implemented in BATIS does not follow the typical, triangle-plot-based strategy. Rather, it sweeps each of the three outer screening gates, i.e., S1, S3, and S4, against all finger gates relevant to the given 1D current channel. The traditional triangle plots [31, 29] are orthogonal to the measured planes shown in Fig. 3(c) and would have been measurements with the S2 axis instead. In other words, the triangle plots and the plots utilized in BATIS contain similar information on different plane cuts through the same space.
The accumulation plots in Fig. 3(c) reveal a dip in the current exiting when the current starts to flow in , shown by the dark green shape imposed on top of the signal in . This is due to a coupling between those two channels resulting from a shared ohmic connection to source the currents. Such a dip in the charge sensor readout is often called a shadow. Shadows increase the difficulty of detecting and forming isolated current channels. The accumulation plot for I1 is much less correlated with the current flow in the two charge sensor channels, but there is still some cross-talk. This is due to separation with the S2 gate which divides the 2DEG between the upper readout channels and the lower dot channel.
Finger Gate Characterization. The last stage of BATIS involves finger gate characterization. This characterization is done individually for all of the finger gates on the QD device. The voltage range for each measurement is always the same and predefined in the configuration file. However, due to the Coulomb blockade oscillations, the automated search for the starting voltage for some gates is more challenging than for others. This can be seen in the raw data pictured in Fig. 3(d).
To ensure that the right point is consistently found, BATIS combines thresholding and slope extraction for each plot to ensure that the pinch-off it finds is robust to noise and spurious fluctuations that do not exceed the threshold. The resulting pinch-off values of all of these varying curves for all finger gates are presented in Table 4. This configuration is used in the final colorized SEM for the ending state of the algorithm shown in Fig. 1(d).
| Gate | VP () | Gate | VP () | Gate | VP () |
| B1 | B6 | B8 | |||
| P1 | P5 | P6 | |||
| B2 | B7 | B9 | |||
| P2 | |||||
| B3 | |||||
| P3 | |||||
| B4 | |||||
| P4 | |||||
| B5 | |||||
IV Summary and outlook
In this paper, we proposed and demonstrated BATIS: a fully automatic procedure for bootstrapping, automated testing, and characterization of accumulation mode, undoped QD devices. The prior knowledge about the QD device required to initialize BATIS is limited to the heterostructure, the device geometry, and the device limits. BATIS navigates through complex high-dimensional gate space to find the final gate configuration in which all desirable current channels are formed following simple physical principles.
The goal of BATIS is to test the device’s functionality to confirm that all gates can properly pinch off and that the current channels can be formed. Testing and characterizing an unknown QD device is essential for identifying suitable global device states that enable the observation of Coulomb blockade. Importantly, the single button press testing and initialization is demonstrated at , eliminating the need for a full cooldown just to learn that a device is not functional. This result is particularly important as we envision a hot large-scale spin-based quantum processor. These systems will require qubit initialization, relying on a complete and accurate device characterization. One could imagine using BATIS to probe the attributes of many devices spread across a wafer or provide device diagnostics on a partition of a large qubit processor without going to temperatures.
The flexible structure allows BATIS to be scalable to characterize and initialize any accumulation mode linear array of QD and easily adaptable to other QD platforms. In particular, BATIS could straightforwardly mimic the behavior of algorithms proposed for nanowires and GaAs/AlxGa1-xAs by simply invoking skip logic on the steps not necessarily supported for these devices. However, it is quite unlikely that any previously proposed algorithms could be easily adapted to the Si/SixGe1-x quad-QD device considered in this work.
Previous studies have demonstrated a device-agnostic algorithm for certain aspects of QD device tuning, such as setting topology [32, 17] and charge state classification [33, 22]. However, all these algorithms presuppose that device characterization and foundational tuning steps—determining gate functionality, defining the QD channel(s), or locating regions of interest for the coarse tuning regime—have already been performed. BATIS fills this gap, making device characterization easily within reach for linear QD array systems and paving the way toward fully automatic control. When linked together, all proposed subroutines can form an autonomous system for controlling arbitrary QD devices up to any desired level of characterization, which will accommodate the growth of QD qubit systems on the path to a scalable quantum computer on this platform.
Acknowledgements.
We acknowledge Danielle Middlebrooks for helping with the automation. We acknowledge HRL Laboratories, LLC for support and L.F. Edge for providing one of the Si/SixGe1-x heterostructures used in this work. This work supported in part by ARO grant nos. W911NF-23-1-0110 and W911NF-17-1-0274. The views and conclusions contained in this paper are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright noted herein. Any mention of commercial products is for information only; it does not imply recommendation or endorsement by NIST.References
- Burkard et al. [2023] G. Burkard, T. D. Ladd, A. Pan, J. M. Nichol, and J. R. Petta, Semiconductor spin qubits, Reviews of Modern Physics 95, 025003 (2023).
- van der Wiel et al. [2002] W. G. van der Wiel, S. De Franceschi, J. M. Elzerman, T. Fujisawa, S. Tarucha, and L. P. Kouwenhoven, Electron transport through double quantum dots, Rev. Mod. Phys. 75, 1 (2002).
- Loss and DiVincenzo [1998] D. Loss and D. P. DiVincenzo, Quantum computation with quantum dots, Phys. Rev. A 57, 120 (1998).
- Hanson et al. [2007] R. Hanson, L. P. Kouwenhoven, J. R. Petta, S. Tarucha, and L. M. K. Vandersypen, Spins in few-electron quantum dots, Rev. Mod. Phys. 79, 1217 (2007).
- Zwanenburg et al. [2013] F. A. Zwanenburg, A. S. Dzurak, A. Morello, M. Y. Simmons, L. C. L. Hollenberg, G. Klimeck, S. Rogge, S. N. Coppersmith, and M. A. Eriksson, Silicon quantum electronics, Rev. Mod. Phys. 85, 961 (2013).
- Ha et al. [2022] W. Ha, S. D. Ha, M. D. Choi, Y. Tang, A. E. Schmitz, M. P. Levendorf, K. Lee, J. M. Chappell, T. S. Adams, D. R. Hulbert, E. Acuna, R. S. Noah, J. W. Matten, M. P. Jura, J. A. Wright, M. T. Rakher, and M. G. Borselli, A flexible design platform for si/sige exchange-only qubits with low disorder, Nano Lett. 22, 1443 (2022), pMID: 34806894.
- Wang et al. [2024] C.-A. Wang, V. John, H. Tidjani, C. X. Yu, A. S. Ivlev, C. Déprez, F. van Riggelen-Doelman, B. D. Woods, N. W. Hendrickx, W. I. L. Lawrie, L. E. A. Stehouwer, S. D. Oosterhout, A. Sammak, M. Friesen, G. Scappucci, S. L. de Snoo, M. Rimbach-Russ, F. Borsoi, and M. Veldhorst, Operating semiconductor quantum processors with hopping spins, Science 385, 447 (2024).
- George et al. [2024] H. C. George, M. T. Mądzik, E. M. Henry, A. J. Wagner, M. M. Islam, F. Borjans, E. J. Connors, J. Corrigan, M. Curry, M. K. Harper, D. Keith, L. Lampert, F. Luthi, F. A. Mohiyaddin, S. Murcia, R. Nair, R. Nahm, A. Nethwewala, S. Neyens, R. D. Raharjo, C. Rogan, R. Savytskyy, T. F. Watson, J. Ziegler, O. K. Zietz, R. Pillarisetty, N. C. Bishop, S. A. Bojarski, J. Roberts, and J. S. Clarke, 12-spin-qubit arrays fabricated on a 300 mm semiconductor manufacturing line (2024), arXiv:2410.16583 [cond-mat.mes-hall] .
- Borsoi et al. [2024] F. Borsoi, N. W. Hendrickx, V. John, M. Meyer, S. Motz, F. Van Riggelen, A. Sammak, S. L. De Snoo, G. Scappucci, and M. Veldhorst, Shared control of a 16 semiconductor quantum dot crossbar array, Nature Nanotechnology 19, 21 (2024).
- Park et al. [2024] S. Park, J. Benson, J. Corrigan, J. P. Dodson, S. N. Coppersmith, M. Friesen, and M. A. Eriksson, Single-shot latched readout of a quantum dot qubit using barrier gate pulsing (2024), arXiv:2408.15380 [cond-mat].
- Vandersypen et al. [2017] L. M. K. Vandersypen, H. Bluhm, J. S. Clarke, A. S. Dzurak, R. Ishihara, A. Morello, D. J. Reilly, L. R. Schreiber, and M. Veldhorst, Interfacing spin qubits in quantum dots and donors—hot, dense, and coherent, npj Quantum Inf. 3, 34 (2017).
- Meirav et al. [1990] U. Meirav, M. A. Kastner, and S. J. Wind, Single-electron charging and periodic conductance resonances in gaas nanostructures, Phys. Rev. Lett. 65, 771 (1990).
- de Kruijf et al. [2023] M. de Kruijf, S. Geyer, T. Berger, M. Mergenthaler, F. Braakman, R. J. Warburton, and A. V. Kuhlmann, A compact and versatile cryogenic probe station for quantum device testing, Review of Scientific Instruments 94, 054707 (2023).
- Neyens et al. [2024] S. Neyens, O. K. Zietz, T. F. Watson, F. Luthi, A. Nethwewala, H. C. George, E. Henry, M. Islam, A. J. Wagner, F. Borjans, E. J. Connors, J. Corrigan, M. J. Curry, D. Keith, R. Kotlyar, L. F. Lampert, M. T. Mądzik, K. Millard, F. A. Mohiyaddin, S. Pellerano, R. Pillarisetty, M. Ramsey, R. Savytskyy, S. Schaal, G. Zheng, J. Ziegler, N. C. Bishop, S. Bojarski, J. Roberts, and J. S. Clarke, Probing single electrons across 300-mm spin qubit wafers, Nature 629, 80 (2024).
- Baart et al. [2016] T. A. Baart, P. T. Eendebak, C. Reichl, W. Wegscheider, and L. M. K. Vandersypen, Computer-automated tuning of semiconductor double quantum dots into the single-electron regime, Appl. Phys. Lett. 108, 213104 (2016).
- Darulová et al. [2020] J. Darulová, S. J. Pauka, N. Wiebe, K. W. Chan, G. C. Gardener, M. J. Manfra, M. C. Cassidy, and M. Troyer, Autonomous tuning and charge-state detection of gate-defined quantum dots, Phys. Rev. Appl. 13, 054005 (2020).
- Zubchenko et al. [2024] A. Zubchenko, D. Middlebrooks, T. Rasmussen, L. Lausen, F. Kuemmeth, A. Chatterjee, and J. P. Zwolak, Autonomous bootstrapping of quantum dot devices, arXiv:2407.20061 (2024).
- Lapointe-Major et al. [2020] M. Lapointe-Major, O. Germain, J. Camirand Lemyre, D. Lachance-Quirion, S. Rochette, F. Camirand Lemyre, and M. Pioro-Ladrière, Algorithm for automated tuning of a quantum dot into the single-electron regime, Phys. Rev. B 102, 085301 (2020).
- Severin et al. [2024] B. Severin, D. T. Lennon, L. C. Camenzind, F. Vigneau, F. Fedele, D. Jirovec, A. Ballabio, D. Chrastina, G. Isella, M. De Kruijf, M. J. Carballido, S. Svab, A. V. Kuhlmann, S. Geyer, F. N. M. Froning, H. Moon, M. A. Osborne, D. Sejdinovic, G. Katsaros, D. M. Zumbühl, G. A. D. Briggs, and N. Ares, Cross-architecture tuning of silicon and SiGe-based quantum devices using machine learning, Sci. Rep. 14, 17281 (2024).
- Zwolak et al. [2020] J. P. Zwolak, T. McJunkin, S. S. Kalantre, J. Dodson, E. R. MacQuarrie, D. Savage, M. Lagally, S. Coppersmith, M. A. Eriksson, and J. M. Taylor, Autotuning of double-dot devices in situ with machine learning, Phys. Rev. Appl. 13, 034075 (2020).
- Zwolak et al. [2021] J. P. Zwolak, T. McJunkin, S. S. Kalantre, S. F. Neyens, E. R. MacQuarrie, M. A. Eriksson, and J. M. Taylor, Ray-based framework for state identification in quantum dot devices, PRX Quantum 2, 020335 (2021).
- Ziegler et al. [2023a] J. Ziegler, F. Luthi, M. Ramsey, F. Borjans, G. Zheng, and J. P. Zwolak, Tuning arrays with rays: Physics-informed tuning of quantum dot charge states, Phys. Rev. Appl. 20, 034067 (2023a).
- Ziegler et al. [2023b] J. Ziegler, F. Luthi, M. Ramsey, F. Borjans, G. Zheng, and J. P. Zwolak, Automated extraction of capacitive coupling for quantum dot systems, Phys. Rev. Appl. 19, 054077 (2023b).
- Rao et al. [2024] A. S. Rao, D. Buterakos, B. van Straaten, V. John, C. X. Yu, S. D. Oosterhout, L. Stehouwer, G. Scappucci, M. Veldhorst, F. Borsoi, and J. P. Zwolak, MAViS: Modular Autonomous Virtualization System for Two-Dimensional Semiconductor Quantum Dot Arrays, arXiv:2411.12516 (2024).
- Zwolak and Taylor [2023] J. P. Zwolak and J. M. Taylor, Colloquium: Advances in automation of quantum dot devices control, Rev. Mod. Phys. 95, 011006 (2023).
- Wolfe et al. [2024] M. A. Wolfe, B. X. Coe, J. S. Edwards, T. J. Kovach, T. McJunkin, B. Harpt, D. Savage, M. G. Lagally, R. McDermott, M. Friesen, S. Kolkowitz, and M. A. Eriksson, Control of threshold voltages in Si/Si0.7 Ge0.3 quantum devices via optical illumination, Physical Review Applied 22, 034044 (2024).
- Zajac et al. [2016] D. M. Zajac, T. M. Hazard, X. Mi, E. Nielsen, and J. R. Petta, Scalable gate architecture for a one-dimensional array of semiconductor spin qubits, Phys. Rev. Appl. 6, 054013 (2016).
- Dodson et al. [2020] J. P. Dodson, N. Holman, B. Thorgrimsson, S. F. Neyens, E. R. MacQuarrie, T. McJunkin, R. H. Foote, L. F. Edge, S. N. Coppersmith, and M. A. Eriksson, Fabrication process and failure analysis for robust quantum dots in silicon, Nanotechnology 31, 505001 (2020).
- Schug et al. [2024] D. Schug, T. J. Kovach, M. A. Wolfe, J. Benson, S. Park, J. P. Dodson, J. Corrigan, M. A. Eriksson, and J. P. Zwolak, Automation of Quantum Dot Measurement Analysis via Explainable Machine Learning (2024), version Number: 4.
- Note [2] The derivative of the screening gate pinch-off was fit to a function like . The non-central screening gates were set to be before the mean of the derivative.
- McJunkin [2021] T. W. McJunkin, Heterostructure Modifications, Fabrication Improvements, and Measurements Automation of Si/SiGe Quantum Dots for Quantum Computation, Ph.D. thesis, The University of Wisconsin-Madison, Madison, WI, USA (2021).
- Ziegler et al. [2022] J. Ziegler, T. McJunkin, E. S. Joseph, S. S. Kalantre, B. Harpt, D. E. Savage, M. G. Lagally, M. A. Eriksson, J. M. Taylor, and J. P. Zwolak, Toward robust autotuning of noisy quantum dot devices, Phys. Rev. Appl. 17, 024069 (2022).
- Durrer et al. [2020] R. Durrer, B. Kratochwil, J. Koski, A. Landig, C. Reichl, W. Wegscheider, T. Ihn, and E. Greplova, Automated tuning of double quantum dots into specific charge states using neural networks, Phys. Rev. Appl. 13, 054019 (2020).